### v2

@@ 2.3.1 Shape Subspace
The similarity between two shape subspaces $\text{sim}(\mathbf{S}_{1},\mathbf{S}_{2})$ can be measured by the canonical angles $\theta_{i}$, where $\{0\leq\theta_{1},\dots,\theta_{N}\leq\frac{\pi}{2}\}$, and $N=3$.
@@
Let $\{\mathbf{\Phi}_{i}\}_{i=1}^{N}$ be the basis vectors of $\mathbf{S}_{1}$ and $\{\mathbf{\Psi}_{i}\}_{i=1}^{N}$ be the basis vectors of $\mathbf{S}_{2}$. The projection matrices $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ are calculated as:
$\mathbf{P}_{1}=\sum_{i=1}^{N}\mathbf{\Phi}_{i}\mathbf{\Phi}_{i}^{T},\quad\mathbf{P}_{2}=\sum_{i=1}^{N}\mathbf{\Psi}_{i}\mathbf{\Psi}_{i}^{T}.$ (7)
$\cos^{2}\theta_{i}$ is the $i$-th largest eigenvalue of $\mathbf{P}_{1}\mathbf{P}_{2}$ or $\mathbf{P}_{2}\mathbf{P}_{1}$.
2.3.2 The Refinement Algorithm
First, landmark coordinate candidates are obtained from heatmaps $f(\mathbf{x})\in\mathbb{R}^{W^{\prime}\times H^{\prime}\times\mathit{K}}$ generated by a model. In each heatmap, up to five candidates for the landmark position are extracted. The candidates are selected in descending order based on their intensity values in the heatmap, ensuring that each candidate is at least a distance of length threshold $\mathit{L}$ away from previously selected candidates to maintain spatial separation. Additionally, the values of the selected candidates must exceed the threshold of $\mathit{r}\%$ of the $\max$ value in the heatmap.
Second, all possible combinations $\mathit{c_{j}}$ of the selected landmark candidates are evaluated based on their similarities to the training data using shape subspace. Here, $\mathit{c_{j}}$ represents the $(x,y)$ coordinates of $\mathit{K}$ landmarks, and $\mathcal{C}$ denotes the set of all such combinations. For each combination $\mathit{c_{j}}\in\mathcal{C}$, its similarity to $\mathit{M}$ training samples is computed, and the mean similarity is taken. Finally, the combination $\mathit{c^{*}}$ that has the highest mean similarity is selected as follows:
@@ 2.3.2 The Refinement Algorithm
2.4 Point-based Bone Segmentation
@@
We demonstrate the segmentation of the humerus and ulna using SAM2, with the detected landmarks as the point prompts.
2.5 Experiments
@@
In this section, we perform landmark detection on our proposed dataset to evaluate the accuracy of joint space measurement. We assessed model performance with different landmark configurations: two-landmark detection (detecting the joint measurement points, landmark 1 and 2) and eight-landmark detection (detecting all landmarks from 1 to 8). The dataset, consisting of 4,201 images from 22 participants, was divided by individuals into 5 groups, with 3 groups used for training, 1 group for validation, and 1 group for testing, with 5-fold cross-validation.
2.5.1 Experimental Configuration
[…]
@@ 2.5.1 Experimental Configuration
    HRNet Three input image sizes $256\times 192$ pixels, $384\times 288$ pixels, $512\times 384$ pixels were compared with corresponding output heatmap sizes of $64\times 48$ pixels, $96\times 72$ pixels, $128\times 96$ pixels, respectively. The augmentation process included random scaling and rotation. The NME was used as the evaluation metric. The total epoch was 210, with the best epoch being selected using the validation set. The optimizer was Adam. The initial learning rate was $1\times 10^{-4}$, and it was adjusted using a step-based policy, with reductions at epochs 170 and 200. The backbone was HRNet-w32. The pre-trained model trained on ImageNet dataset was used. The joint weights of all the landmarks were set to 1.0.
[…]
@@
    U-Net Two input image sizes $256\times 256$ pixels and $512\times 512$ pixels were used, with output heatmap sizes matching the input sizes. The augmentation process included random scaling and rotation. The normalized mean error was used as the evaluation metric. The total epoch was 600, with the best epoch selected by the validation set, and the learning rate was set to $1\times 10^{-3}$.
[…]
@@
    PCT Two input image sizes were evaluated: $256\times 256$ pixels. Data augmentation included random scaling, rotation, and shift. The Swin Transformer V2 backbone pre-trained with SimMIM on ImageNet was used. The optimizer was AdamW, with an initial learning rate of $8\times 10^{-4}$, which was adjusted using a cosine annealing schedule. The number of tokens was set to 34. The training epoch was set to 50 for the tokenizer and 210 for the classifier. The joint weights of all the landmarks were set to 1.0.
[…]
@@
    YOLOv8 The input image size was $640\times 640$ pixels. The augmentation process included random rotations, scaling, shifts, and horizontal flips. Training was conducted for 1,000 epochs, with the best epoch selected for evaluation. The optimizer was AdamW for the first $10,000$ iterations, followed by SGD. The initial learning rate was $1\times 10^{-3}$ for AdamW and $1\times 10^{-2}$ for SGD. The pose loss weight was set to 12.0.
  • The Number of Landmarks
@@
    We compared two-landmark and eight-landmark detection. The two-landmark detection model detects landmarks 1 and 2. The eight-landmark model detects all landmarks from 1 to 8. The landmark refinement is only applied to eight-landmark detection.
  • The Landmark Refinement
@@
    The value threshold $r$ was set to 75%, and the length threshold $L$ was set to 5 pixels on the heatmap for the candidate selection. We used random $0.033\%$ of the total training samples for the similarity calculation, which, for instance, corresponds to $M=67$.
  • The Point Prompts for Point-based Segmentation
@@
    We used the detected 6 points in the eight-landmark detection as the positive point prompts $(x_{i},y_{i}),\ (i=3,..,8)$. Also, we set the 6 negative point prompts based on the detected landmark coordinates $(x_{4},y_{3}),(x_{4},y_{3}),\ (x_{5},y_{4}),\ (x_{5},2y_{5}-y_{4}),\ (x_{8},2y_{5}-y_{4}),\ ((x_{1}+x_{2})/2,(y_{1}+y_{2})/2)$.
2.5.2 Evaluation Metrics
We used Mean Absolute Error (MAE) as a metric to evaluate the accuracy of each predicted landmark. Also, Euclidean Distance Error (EDE) is used to evaluate the joint space length. For metric definitions, please see the Supplementary Information.
$\text{MAE}_{k}=\frac{1}{N}\sum_{i=1}^{N}{\sqrt{(x_{i,k}-x_{i,k}^{\text{gt}})^{2}+(y_{i,k}-y_{i,k}^{\text{gt}})^{2}}}.$ (9)
where $N$ represents the total number of test samples, $x_{j,k}$ and $y_{j,k}$ are the predicted $x$- and $y$-coordinates of the $k$-th landmark in the $i$-th image, and $x_{i,k}^{\text{gt}}$ and $y_{i,k}^{\text{gt}}$ are the ground truth coordinates of the same landmark.
Also, Euclidean Distance Error (EDE) is used to evaluate the joint space length. The EDE is defined as the mean Euclidean distance error across all $N$ samples.
$\text{EDE}=\frac{1}{N}\sum_{i=1}^{N}\left|{\|\mathbf{p_{1,i}}-\mathbf{p_{1,i}}\|}_{2}-{\|\mathbf{p_{2,i}^{\text{gt}}}-\mathbf{p_{2,i}^{\text{gt}}}\|}_{2}\right|.$
where ${\|\mathbf{p_{1,i}}-\mathbf{p_{1,i}}\|}_{2}$, ${\|\mathbf{p_{2,i}^{\text{gt}}}-\mathbf{p_{2,i}^{\text{gt}}}\|}_{2}$ represent the Euclidean distances for the predicted and ground truth lengths between landmarks 1 and 2, respectively.
3 Results
@@
Table 1: The two-landmark detection results. LM1 and LM2 represent the Mean Absolute Error (MAE) (mm) for landmarks 1 and 2, respectively. Ave2 is the average MAE of two landmarks. Length denotes the mean Euclidean distance error (mm) for joint space length. The bold values indicate the lowest error in each column. Smaller values indicate better detection performance.
@@
ViTPose 256×192 0.195 0.233 0.214 0.249
384×288 0.180 0.307 0.243 0.293
512×384 0.200 0.198 0.199 0.210
HRNet 256×192 0.187 0.227 0.207 0.221
384×288 0.149 0.137 0.143 0.140
512×384 0.125 0.118 0.122 0.116
U-Net 256×256 0.230 0.188 0.209 0.145
512×512 0.174 0.150 0.162 0.112
YOLOv8 640×640 1.878 2.526 2.202 1.027
PCT 256×256 0.552 0.962 0.757 0.803
[…]
Figure 4: The two-landmark detection results with enlarged views around the joints. The upper row displays the ground truth and the lower row illustrates the inference results of HRNet with the input image size $512\times 384$. The pink dot is landmark1, and the light blue dot is landmark2.
Table 2: The eight-landmark detection results. LM1-LM8 represent Mean Absolute Error (MAE) (mm) for landmarks 1 through 8. Ave8 is the average MAE for all 8 landmarks (LM1–LM8), and Ave2 is the average MAE for LM1 and LM2. Length denotes the mean Euclidean distance error (mm) for joint space. SS is Shape Subspace. The bold values indicate the lowest error in each column. Smaller values indicate better detection performance.
@@
ViTPose 256×192 0.210 0.251 0.424 0.463 0.682 0.213 0.245 0.468 0.369 0.231 0.260
384×288 0.183 0.268 0.604 0.741 1.072 0.186 0.266 0.784 0.513 0.225 0.249
512×384 0.237 0.344 1.205 1.164 2.147 0.237 0.323 1.003 0.833 0.291 0.322
ViTPose+SS 256×192 0.210 0.251 0.427 0.463 0.674 0.213 0.245 0.468 0.369 0.231 0.260
384×288 0.183 0.259 0.566 0.608 0.798 0.186 0.258 0.594 0.432 0.221 0.245
512×384 0.201 0.303 0.881 0.971 1.315 0.218 0.262 0.697 0.606 0.252 0.285
HRNet 256×192 0.184 0.251 0.458 0.453 0.774 0.184 0.245 0.495 0.381 0.218 0.239
384×288 0.171 0.183 0.411 0.459 0.781 0.171 0.175 0.472 0.353 0.177 0.185
512×384 0.148 0.154 0.495 0.529 0.976 0.152 0.147 0.471 0.384 0.151 0.158
HRNet+SS 256×192 0.184 0.251 0.458 0.453 0.774 0.184 0.245 0.495 0.381 0.218 0.239
384×288 0.171 0.182 0.383 0.459 0.771 0.171 0.174 0.472 0.348 0.176 0.183
512×384 0.148 0.154 0.419 0.528 0.844 0.152 0.148 0.468 0.358 0.151 0.158
U-Net 256×256 0.202 0.164 0.691 1.000 1.163 0.207 0.163 0.482 0.509 0.183 0.141
512×512 0.175 0.173 5.556 1.666 3.326 0.180 0.148 0.698 1.490 0.174 0.150
YOLOv8 640×640 0.585 0.717 0.677 0.540 0.717 0.587 0.710 1.837 0.796 0.651 0.718
PCT 256×256 0.640 0.954 1.087 0.655 1.235 0.623 0.941 1.035 0.896 0.797 0.585
@@
3.1 The Two-Landmark Detection
Table 1 displays the results for the two-landmark detection. As shown in the table, the heatmap-based methods outperformed other approaches. Among the heatmap-based methods, HRNet with an input image size of $512\times 384$ pixles, achieved the best performance. The average MAE for the two landmarks was $0.122$, and the EDE for the joint space was $0.116$ in this configuration. Fig. 4 shows the detected landmarks. As illustrated in this figure, the positions of landmarks 1 and 2 are precisely detected, even in the presence of variations in bone shapes and white noise in the joint space.
Figure 5: The eight-landmark detection results with enlarged views around the joints. The first row shows the ground truth for 8 landmarks. The second row illustrates the positions of detected landmarks by HRNet+SS with the input size $384\times 288$. The colors of landmarks 1 to 8 are pink, light blue, red, yellow, green, blue, orange, and purple.
Figure 6: The point-based segmentation results using SAM2 with the point prompts calculated by the detected landmark positions. (a) The original images, (b) The ground truth of the bone segmentation. (c) The segmented images using SAM2 and the point prompts calculated from detected landmarks. The blue areas represent the segmented regions, the green stars indicate the positive point prompts and the red stars denote the negative point prompts.
Figure 7: An example of landmark refinement using shape subspace. (a) shows overlays of the output heatmaps on the input image. In the LM3 heatmap, multiple peaks are observed, with the right peak exhibiting higher intensity. However, shape subspace refinement correctly selects the left peak, which more accurately aligns with the ground truth. (b) displays the detected landmarks, where the red point represents the initial position obtained from LM3. The detected landmarks 1–8 are shown in pink, light blue, red, yellow, green, blue, orange, and purple, respectively.
3.2 The Eight-Landmark Detection

### v3

@@ IV-B1 Shape Subspace
The similarity between two shape subspaces $\text{sim}(\mathbf{S}_{1},\mathbf{S}_{2})$ can be measured by the canonical angles $\theta_{i}$, where $\{0\leq\theta_{1},\dots,\theta_{N}\leq\frac{\pi}{2}\}$, with $N$ representing the subspace dimension. In this case, $N=3$.
@@
Let $\{\mathbf{\Phi}_{i}\}_{i=1}^{N}$ be the basis vectors of $\mathbf{S}_{1}$ and $\{\mathbf{\Psi}_{i}\}_{i=1}^{N}$ be the basis vectors of $\mathbf{S}_{2}$. The projection matrices $\mathbf{Q}_{1}$ and $\mathbf{Q}_{2}$ are calculated as:
$\mathbf{Q}_{1}=\sum_{i=1}^{N}\mathbf{\Phi}_{i}\mathbf{\Phi}_{i}^{T},\quad\mathbf{Q}_{2}=\sum_{i=1}^{N}\mathbf{\Psi}_{i}\mathbf{\Psi}_{i}^{T}.$ (7)
$\cos^{2}\theta_{i}$ is the $i$-th largest eigenvalue of $\mathbf{Q}_{1}\mathbf{Q}_{2}$ or $\mathbf{Q}_{2}\mathbf{Q}_{1}$ (Algorithm 2).
Input: Shape matrix $X\in\mathbb{R}^{K\times N}$
Output: Basis vectors $S\in\mathbb{R}^{K\times N}$ of shape subspace
$\mu\leftarrow$ column-wise mean of $X$;
$X_{c}\leftarrow X-\mathbf{1}\mu^{\top}$;
$U,\Sigma,V^{T}]\leftarrow\mathrm{SVD}(X_{c})$;
$S\leftarrow U[:,:N]$;
return $S$
Algorithm 1 ShapeSubspaceBasis($X$, $N=3$)
Input: Subspace basis $\Phi\in\mathbb{R}^{K\times N}$ for $S_{1}$, subspace basis $\Psi\in\mathbb{R}^{K\times N}$ for $S_{2}$
Output: Similarity $\text{sim}(S_{1},S_{2})\in\mathbb{R}$
$Q_{1}\leftarrow\Phi\Phi^{\top}$;
$Q_{2}\leftarrow\Psi\Psi^{\top}$;
$\text{sim}\leftarrow\frac{1}{N}\mathrm{trace}(Q_{1}Q_{2})$;
return sim
Algorithm 2 ShapeSubspaceSimilarity($\Phi$, $\Psi$, $N=3$)
Figure 4: The two-landmark detection results shown in enlarged views around the joints. The upper row displays the ground truth, and the lower row illustrates the inference results of HRNet. The pink dot is landmark1, and the light blue dot is landmark2.
IV-B2 Refinement Algorithm
All possible combinations $\mathit{c_{j}}$ of the selected landmark candidates are evaluated based on their similarities to the training data using shape subspace. Here, $\mathit{c_{j}}$ represents a combination of $(x,y)$ coordinates of $\mathit{N_{K}}$ landmarks (i.e., the shape matrix $X_{\text{test},\,c_{j}}$), and $\mathcal{C}$ denotes the set of all such combinations. For each combination $\mathit{c_{j}}\in\mathcal{C}$, its similarity to $\mathit{M}$ training samples is computed, and the mean similarity is taken. Finally, the combination $\mathit{c^{*}}$ that has the highest mean similarity is selected as follows:
@@ IV-B2 Refinement Algorithm
IV-C Point-based Bone Segmentation
@@
The segmentation of the humerus and ulna using SAM2 is demonstrated using the detected landmarks as the point prompts.
V Experiments
@@
In this section, landmark detection was performed on our proposed dataset to evaluate the accuracy of joint space measurement. Model performance was assessed with two different landmark configurations: two-landmark detection (detecting the joint measurement points, landmarks 1 and 2) and eight-landmark detection (detecting all landmarks from 1 to 8).
V-A Experimental Configuration
V-A1 Landmark Detection Models
  • ViTPose
@@ V-A1 Landmark Detection Models
    The input image size was $512\times 384$ pixels, and the output heatmap size was $128\times 96$ pixels. The augmentation process included random scaling and rotation. The Normalized Mean Error (NME) was used as the evaluation metric. The total epoch was 210, with the best epoch being selected using the validation set. The optimizer was Adam. The initial learning rate was $1\times 10^{-4}$, and it was adjusted using a step-based policy, with reductions at epochs 170 and 200. The backbone was ViT-small, and the head was the classic decoder. The pre-trained model trained on ImageNet dataset was used. The patch size was $16\times 16$ pixels. The joint weights of all the landmarks were set to 1.0.
  • HRNet
@@
    The input image size was set to $512\times 384$ pixels, and the output heatmap size was set to $128\times 96$ pixels. The augmentation process included random scaling and rotation. The NME was used as the evaluation metric. The total epoch was 210, with the best epoch being selected using the validation set. The optimizer was Adam. The initial learning rate was $1\times 10^{-4}$, and it was adjusted using a step-based policy, with reductions at epochs 170 and 200. The backbone was HRNet-w32. The pre-trained model trained on ImageNet dataset was used. The joint weights of all the landmarks were set to 1.0.
  • U-Net
    The input image size and the output heatmap size were set to $512\times 512$ pixels. The augmentation process included random scaling and rotation. The normalized mean error was used as the evaluation metric. The total epoch was 600, with the best epoch selected by the validation set, and the learning rate was set to $1\times 10^{-3}$.
  • PCT
@@
    The input image size was set to $512\times 512$ pixels. Data augmentation included random scaling, rotation, and shift. The Swin Transformer V2 backbone pre-trained with SimMIM on ImageNet was used. The optimizer was AdamW, with an initial learning rate of $8\times 10^{-4}$, which was adjusted using a cosine annealing schedule. The number of tokens was set to 34. The training epoch was set to 50 for the tokenizer and 210 for the classifier. The joint weights of all the landmarks were set to 1.0. These parameters were set following the official implementation [30].
  • YOLOv8
@@
    The input image size was $640\times 640$ pixels. The augmentation process included random rotations, scaling, shifts, and horizontal flips. Training was conducted for 1,000 epochs, with the best epoch selected for evaluation. The optimizer was AdamW for the first $10,000$ iterations, followed by SGD. The initial learning rate was $1\times 10^{-3}$ for AdamW and $1\times 10^{-2}$ for SGD. The pose loss weight was set to 12.0.
Figure 5: The eight-landmark detection results. The first row shows the ground truth for 8 landmarks. The second row illustrates the positions of detected landmarks by HRNet+SSR. The colors of landmarks 1 to 8 are pink, light blue, red, yellow, green, blue, orange, and purple.
V-A2 Number of Landmarks
@@
Two-landmark and eight-landmark detection are compared. The two-landmark detection model detects landmarks 1 and 2. The eight-landmark model detects all landmarks from 1 to 8. The landmark refinement is only applied to the eight-landmark detection.
V-A3 Parameter Settings for SSR
@@
For the landmark refinement, the value threshold $r$ was set to 75%, and the length threshold $D$ was set to 5 pixels on the heatmap for the candidate selection. To calculate the similarity, random $3.3\%$ of the total training samples was used, which, for instance, corresponds to $M=67$.
V-A4 Point Prompts for the Point-based Segmentation
@@
We used the detected 6 points in the eight-landmark detection as the positive point prompts $(x_{i},y_{i}),\ (i=3,..,8)$. Also, six negative point prompts based on the detected landmark coordinates $(x_{4},y_{3}),(x_{4},y_{3}),\ (x_{5},y_{4}),\ (x_{5},2y_{5}-y_{4}),\ (x_{8},2y_{5}-y_{4}),\ ((x_{1}+x_{2})/2,\ (y_{1}+y_{2})/2)$ are used.
@@
TABLE I: The two-landmark detection results. LM1 and LM2 represent the Mean Absolute Error (MAE) ($m$) for landmarks 1 and 2, respectively. Ave2 is the average MAE of the two landmarks. Length denotes the mean Euclidean distance error ($m$) for joint space length. The bold values indicate the lowest error in each column. Smaller values indicate better detection performance.
@@
YOLOv8 0.793 1.753 1.340 1.276
PCT 0.579 0.762 0.848 0.730
U-Net 0.239 0.150 0.194 0.212
ViTPose 0.168 0.161 0.164 0.172
HRNet 0.132 0.129 0.131 0.125
TABLE II: The eight-landmark detection results. LM1-LM8 represent Mean Absolute Error (MAE) ($m$) for landmarks 1 through 8. Ave8 is the average MAE for all 8 landmarks (LM1–LM8), and Ave2 is the average MAE for LM1 and LM2. Length denotes the mean Euclidean distance error ($m$) for joint space. SSR is Shape Subspace Refinement. The bold values indicate the lowest error in each column. Smaller values indicate better detection performance.
@@
YOLOv8 0.568 0.664 0.949 0.584 0.871 0.571 0.646 1.127 0.706 0.616 0.574
PCT 0.624 0.833 0.919 0.711 1.301 0.637 0.841 0.903 0.860 0.728 0.708
U-Net 0.204 0.146 3.252 2.729 2.521 0.234 0.137 0.657 1.235 0.175 0.175
U-Net + SSR 0.189 0.131 2.532 2.514 2.346 0.209 0.130 0.539 1.074 0.160 0.159
ViTPose 0.188 0.248 0.775 0.864 1.601 0.193 0.240 0.695 0.600 0.218 0.235
ViTPose + SSR 0.170 0.227 0.612 0.757 1.030 0.183 0.207 0.526 0.464 0.199 0.217
HRNet 0.152 0.175 0.571 0.548 1.018 0.158 0.170 0.443 0.404 0.164 0.169
HRNet + SSR 0.152 0.175 0.515 0.536 0.891 0.158 0.169 0.442 0.380 0.163 0.169
TABLE III: Fold-wise two- and eight-landmark detection results. Ave2 is the average MAE for LM1 and LM2, and Ave8 is the average MAE for all 8 landmarks (LM1–LM8). Length denotes the mean Euclidean distance error for joint space. SSR is Shape Subspace Refinement. All values are shown as mean ± standard deviation (in $m$). Smaller values indicate better detection performance.
Two-landmark detection Eight-landmark detection
HRNet HRNet+SSR
@@
fold 1-1 0.141 ± 0.141 0.153 ± 0.182 0.460 ± 1.118 0.196 ± 0.129 0.269 ± 0.228
fold 1-2 0.147 ± 0.106 0.106 ± 0.148 0.338 ± 0.383 0.188 ± 0.115 0.127 ± 0.158
fold 1-3 0.158 ± 0.308 0.195 ± 0.442 0.420 ± 0.686 0.188 ± 0.150 0.208 ± 0.263
fold 1-4 0.125 ± 0.077 0.076 ± 0.092 0.429 ± 1.125 0.140 ± 0.083 0.136 ± 0.136
fold 1-5 0.130 ± 0.091 0.127 ± 0.143 0.361 ± 0.623 0.165 ± 0.091 0.155 ± 0.193
fold 2-1 0.118 ± 0.081 0.127 ± 0.159 0.341 ± 0.528 0.150 ± 0.150 0.169 ± 0.176
fold 2-2 0.125 ± 0.141 0.113 ± 0.183 0.421 ± 1.382 0.158 ± 0.158 0.166 ± 0.436
fold 2-3 0.127 ± 0.478 0.130 ± 0.426 0.286 ± 0.276 0.160 ± 0.160 0.182 ± 0.218
fold 2-4 0.128 ± 0.078 0.138 ± 0.166 0.415 ± 0.690 0.162 ± 0.162 0.175 ± 0.208
fold 2-5 0.111 ± 0.081 0.083 ± 0.106 0.326 ± 0.395 0.126 ± 0.126 0.101 ± 0.121
V-B Evaluation Metrics
Mean Absolute Error (MAE) was used to evaluate the accuracy of each predicted landmark.
$\text{MAE}_{k}=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}{\sqrt{(x_{i,k}-x_{i,k}^{\text{gt}})^{2}+(y_{i,k}-y_{i,k}^{\text{gt}})^{2}}}.$ (9)
where $N_{t}$ represents the total number of test samples, $x_{i,k}$ and $y_{i,k}$ are the predicted $x$- and $y$-coordinates of the $k$-th landmark in the $i$-th image, and $x_{i,k}^{\text{gt}}$ and $y_{i,k}^{\text{gt}}$ are the ground truth coordinates of the same landmark.
Also, Euclidean Distance Error (EDE) was used to evaluate the joint space length. The EDE is defined as the mean Euclidean distance error across all $N_{t}$ samples.
$\text{EDE}=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\left|{\|\mathbf{p_{1,i}}-\mathbf{p_{2,i}}\|}_{2}-{\|\mathbf{p_{1,i}^{\text{gt}}}-\mathbf{p_{2,i}^{\text{gt}}}\|}_{2}\right|.$
where ${\|\mathbf{p_{1,i}}-\mathbf{p_{2,i}}\|}_{2}$, ${\|\mathbf{p_{1,i}^{\text{gt}}}-\mathbf{p_{2,i}^{\text{gt}}}\|}_{2}$ represent the Euclidean distances for the predicted and ground truth lengths between landmarks 1 and 2, respectively.
Our dataset of 22 participants was split into 5 groups based on individuals. We performed 5-fold cross-validation under two different split settings. In each fold, 3 groups were used for training, 1 for validation, and 1 for testing; for example, one such split included 13 participants for training, 4 for validation, and 5 for testing. The full-size images without cropping were used in the experiment. The final results were averaged over all ten folds across the two split settings. The samples included in each split are detailed in the supplementary material.
VI Results
VI-A The Two-Landmark Detection
Table I presents the results of two-landmark detection, showing that heatmap-based methods outperformed other approaches. Among them, HRNet achieved the best performance, with an average MAE of $0.131$ for the two landmarks and an EDE of $0.125$ for the joint space. As illustrated in Fig. 4, the positions of landmarks 1 and 2 are precisely detected, even in the presence of variations in bone shapes and white noise in the joint space. Fold-wise results for HRNet are detailed in Table III. To further demonstrate the applicability of our method to diagnose UCL injury, joint space measurements were conducted with and without 2.5 kgf of valgus stress applied proximally to the wrist. Fig. 8 shows HRNet-based inference results on the medial elbow of a baseball player under both stressed and unstressed conditions, demonstrating accurate detection in both cases. Additional results from eight baseball players not included in the dataset are provided in the supplementary material.
VI-B The Eight-Landmark Detection